# Gravitational wave probes of particle dark matter: a review

### v3

@@ IX.2 Dephasing due to dynamical friction of IMRI systems
PBHs could also host DM spikes, and must do so, since, at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al. (2024). Thus, Ref. Cole et al. (2023a) showed that ET and Cosmic Explorer could probe the existence DM spikes around binary PBH systems, where the primary has a mass of $[1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$. However, just as in astrophysical binary black hole systems, these exotic mergers would be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer. Multi-band observation of inspirals/mergers constraining DM spikes are also possible with future deci-hertz and ground-based GW interferometers, which would enhance the constraints on the power-law index and spike-density factors by orders of magnitude relative to individual constraints from each of these detectors Tahelyani et al. (2024); Chen et al. (2024b).
Recently, it has also been shown that these dephasing effects are unique enough to allow distinguishing between three BH environments: accretion disks, DM spikes, and ultralight boson clouds Cole et al. (2023b), as well as distinguishing between environmental effects and modified theories of gravity Yuan et al. (2024). However, Ref. Zwick et al. (2024) argues that practically speaking, dephasing could result from a variety of (also unmodelled) sources, and would be difficult (computationally and algorithmically speaking) to track with long-duration templates in matched filtering. Though, Ref. Wilcox et al. (2024) has indicated this could be possible within a parameterized post-Einsteinian framework. They thus propose more model-agnostic ways (i.e. periodic forces) that may influence the evolution of the binary and leave characteristic signatures on the GW signal. Such problems with matched filter, could, however, be handled by the use of alternative, non-matched filtering CW methods meant to handle intermediate-duration signals, e.g. Isi et al. (2019); Sun and Melatos (2019); Miller et al. (2018); Oliver et al. (2019); Banagiri et al. (2019); Miller et al. (2019); Guo and Miller (2022); Andrés-Carcasona et al. (2023); Alestas et al. (2024); Miller et al. (2024a); Andrés-Carcasona et al. (2025), and those discussed already in this review.
Perhaps more surprising, it has been proposed that equal-mass systems could sustain DM spikes even in the last ten orbits of binary black hole merger Aurrekoetxea et al. (2024b). Initially, it had been shown through N-body simulations that heavy, particle-like DM environments would be disrupted well before equal-mass binary black hole systems merge Bertone and Merritt (2005); Kavanagh et al. (2018). But, for wave-like, ultralight dark matter, though there have been many works that have modeled the motion of binary black holes through DM environments Koo et al. (2024); Boudon et al. (2023); Zhong et al. (2023); Cardoso et al. (2022a, b); Vicente and Cardoso (2022); Annulli et al. (2020a, b); Brax et al. (2020); Bromley et al. (2024) or with boson clouds Cao and Tang (2023); Traykova et al. (2023); Chia et al. (2023); Baumann et al. (2019a, 2020); Ikeda et al. (2021); Cardoso et al. (2020); Hannuksela et al. (2019); Baumann et al. (2022b); Zhang and Yang (2020); Kumar Poddar et al. (2020); Wong (2019, 2020); Kavic et al. (2020); East (2018); Siemonsen and East (2020); Baumann et al. (2022a); Siemonsen et al. (2023b); Tomaselli et al. (2023); Poddar et al. (2024); Cao et al. (2024); Guo et al. (2024d), the point during the inspiral at which the environment becomes disrupted is not clear, motivating further studies into this Bamber et al. (2023). Ref. Aurrekoetxea et al. (2024b) found that amount of dephasing that exists within the last ten orbits of a binary black hole merger depends strongly the scalar DM mass and the masses of the BHs:
$\mu\approx 5\times 10^{-17}\,\left(M/10^{6}M_{\odot}\right)^{-1}\,\mathrm{eV},$ (109)
which gives the optimal DM mass to cause the most dephasing during the last ten orbits of the binary, and determines the DM density as
$\rho_{0}\approx 10^{20}\,\left(\frac{M}{10^{6}M_{\odot}}\right)^{-2}\,\frac{M_{\odot}}{\mathrm{pc}^{3}}\,.$ (110)
which gives the DM density necessary to induce dephasing. Though the DM density presented here is larger than the average, it results from a measured dephasing over a short period of time. If the effect of the DM spike on the binary black hole system turns on at larger separations, as may be expected from previous works Cole et al. (2023a, b), then smaller densities could cause significant dephasing over the lifetime of the binary.
IX.3 Upper limits on BH environments
@@
Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals. Knowing this, Ref. Caneva Santoro et al. (2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small with respect to the phase evolution of the binary black hole systems in vacuum. While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment. Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3 Gondolo and Silk (1999); Sadeghian et al. (2013). The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger. These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso and Maselli (2020); Kavanagh et al. (2020).
@@ X GW probes of atomic DM
X.1 Background
It is possible that the dark sector is actually much more complicated than a single new particle with a single coupling to gravity or the standard model. In fact, a atomic DM models exist Feng et al. (2009); Kaplan et al. (2010) composed of dark protons and dark electrons that interact via a massless dark photon with a particular coupling. With such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisional excitation processes Buckley and DiFranzo (2018). If DM collects in small-enough regions of space, it will collapse and form “dark BHs Shandera et al. (2018).
Such dark BHs could form in binary systems, as astrophysical BHs do; thus, LIGO, Virgo and KAGRA observations of binary black hole mergers can constrain the existence of these objects, just as they do for PBHs Singh et al. (2021); Abbott et al. (2022f, 2023).
X.2 Constraints from observations of binary black holes
Ref. Shandera et al. (2018) first considered the prospects of detecting dark BH mergers with advanced LIGO and ET, finding that these interferometers could set stringent constraints on the fraction of DM that dark BHs could compose. If the BH merger of GW190425 Abbott et al. (2020a) is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar (1931) can be constrained to be below $1.4M_{\odot}$ at $>99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh et al. (2021).
These constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar (1931); Shandera et al. (2018); Singh et al. (2021):
@@ X.2 Constraints from observations of binary black holes
The range of the dark proton mass $m_{\chi}$ can be obtained by setting $M_{\rm DC}=M_{\rm min}$ for the values of $M_{\rm min}$ for which the fraction of DM that dark BHs could compose is less than 1:
$\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm max}}}m_{p}<m_{\chi}<\sqrt{\frac{1.4M_{\odot}}{M_{\rm max}^{\rm max}}}m_{p}$ (112)
$M_{\rm min}$ is the minimum mass in a distribution of dark BHs masses that is assumed, as described below.
@@
Further constraints on the fraction of DM composed of dark BHs are presented in three cases, assuming that (1) none of the detected BHs by LIGO, Virgo and KAGRA are dark, (2) half are dark, and (3) GW190425 and/or GW190814 Abbott et al. (2020b) are dark. Here, we limit ourselves to case (1) and case (3). These constraints rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm and Larson (2004), which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $[M_{\rm min},rM_{\rm min}]$. The parameters $b=[-1,2],r=[2,100]$ determine the mass distribution of dark BHs. With these assumptions, we show in Fig. 27 the constraint on the fraction of DM that dark BHs could compose as a function of the average mass of the BHs considered, in relation to other limits, in case (1) and in case (3). Assuming all BHs are not dark, the constraint beats existing ones by several orders of magnitude; assuming that GW190425 and GW190814 are dark BHs, a region in this parameter space is carved out of possible average mass/ DM fractions. More observations of dark BHs would be needed to further narrow this region. However, it must be emphasized that while GW190425 and GW190814 being dark BHs are not yet excluded by other experiments, this work does not provide a way of distinguishing astrophysical BHs from dark BHs.
@@
Figure 27: Taken from Singh et al. (2021). Constraints on the fraction of DM that dark BHs could compose relative to other experimental upper limits using all binary black holes detected in the first three observing runs of LIGO, Virgo and KAGRA. These plots assume an atomic model of DM that forms dark BHs. Even in the absence of any dark BHs, the constraints obtained in Singh et al. (2021) significantly outperform existing microlensing experiments Allsman et al. (2001); Tisserand et al. (2007); Niikura et al. (2019) and those from dynamics of dwarf galaxies Brandt (2016); Koushiappas and Loeb (2017). Additional constraints arise from PBH searches Abbott et al. (2019); Gaggero et al. (2017); Raidal et al. (2017, 2019); Ali-Haimoud et al. (2017). SB stands for stochastic background and refers to analyses of the first and second observing runs of LIGO, Virgo and KAGRA data for stochastic GW backgrounds from generic systems.
@@
X.3 Constraints from sub-solar mass binary black hole searches
@@
The lack of smoking-gun evidence for dark BHs in current LIGO, Virgo and KAGRA observations motivates the need to consider sub-solar mass BHs, whose origins, if detected, cannot be explained by astrophysical means. Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $[10^{-11},1]M_{\odot}$ Nitz and Wang (2021); Phukon et al. (2021); Abbott et al. (2022f); Miller et al. (2022b); Nitz and Wang (2022); Abbott et al. (2023, 2022b); Miller et al. (2024b), and, traditionally, have put constraints on the fraction of DM that PBHs could compose. However, as discussed in Singh et al. (2021), these searches do not assume a particular formation mechanism for these compact objects, and the binaries can thus be interpreted as arising from not only PBHs, but also dark BHs.
@@
In the most recent analysis of LIGO, Virgo and KAGRA data from the second half of the third observing run, constraints on dark BHs with masses below $1M_{\odot}$ were placed, as shown in Fig. 28, using different matched filtering methods, again assuming a power-law dependence for the mass distribution of dark BHs and as a function of the minimum mass of this distribution. Nuisance parameters in the mass distribution are marginalized over. These limits do not go below $\sim 10^{-2}M_{\odot}$ because the maximum mass probed is between $[2,1000]M_{\odot}$, and so the distribution tends to disfavor smaller masses. Moreover, the limits do not go above a solar mass because the search is only for sub-solar mass objects.
@@
From this work Abbott et al. (2023), the mass of the dark proton has been constrained to be [0.66,8.8] GeV, which is actually broader than the constraint in Singh et al. (2021) of [0.91,4.76] GeV since there are no assumed detections of dark BHs. Additionally, the energy spacing between dark atomic energy levels can also be constrained, assuming values for the mass of the dark fermions, the dark fine structure constant, and the Jeans mass, and using Eq. 1 of Singh et al. (2021) and Eq. 4 of Shandera et al. (2018).
@@
Figure 28: Taken from Abbott et al. (2023). Constraints on the fraction of DM that dark BHs could compose in the atomic DM scenario. A mass distribution for dark BHs is assumed with the shown minimum mass on the $x$-axis, and the constraints are shown for three different matched filtering algorithms: GstLAL (dotted), MBTA (dashed) and PyCBC (solid). Constraints from a previous search for sub-solar mass compact objects in the first half of the third observing run of LIGO, Virgo and KAGRA are also shown  Abbott et al. (2022f).

### v4

@@ IX.2 Dephasing due to dynamical friction of inspiraling primordial black holes
PBHs could also host DM spikes, since at most masses, the fraction of DM that they could compose is less than one Green and Kavanagh (2021); Carr et al. (2024). Thus, Ref. Cole et al. (2023a) showed that ET and Cosmic Explorer could probe the existence of DM spikes around binary PBH systems, where the primary has a mass of $1,10]M_{\odot}$ and the system has a mass ratio of $10^{-3}$. However, similarly to IMRI systems discussed in the previous subsection, these exotic mergers could be missed if the DM spike is not taken into account in matched-filtering searches in ET and Cosmic Explorer. As mentioned in Sections IX.1 and VII.2, one possible approach is to adapt CW methods to search for DM spikes around binary black holes. This could provide a quick, though less sensitive, way to obtain point estimates for important binary black hole parameters, such as the chirp mass. These estimates could then be used to reduce the number of templates needed to account for DM spikes, enabling a follow-up matched filtering search.
IX.3 Upper limits on black-hole environments
@@
Because vacuum phase-evolution waveforms confirmed the detection of many binary black hole mergers by LIGO, Virgo and KAGRA, any deviations due to environmental effects, such as a DM spike, must be small enough to be inconsequential to the detected GW signals. Knowing this, Ref. Caneva Santoro et al. (2024) modeled dynamical friction and other possible environmental effects as additional (negative) PN corrections to the waveform, and placed upper limits on the density of environment around BHs by requiring the induced dephasing to be small enough to not alter the phase evolution of the binary black hole systems in vacuum. While the results were used to eliminate dynamical fragmentation Loeb (2016) as a possible binary formation channel, they are generic enough to apply to any BH environment. Unfortunately, the upper limits on environments affected by dynamical friction, which would be relevant for the DM spike scenario, are, at best, about seven orders of magnitude larger than the maximum density of the spike ($10^{-6}$ g/cm3) Gondolo and Silk (1999); Sadeghian et al. (2013). The upper limits are shown to drastically improve in DECIGO because the dynamical friction (and other environments) affect the early inspiral of the binary black hole system much more than at the time of merger. These findings are consistent with those found in N-Body simulations about how the dynamics of the DM spike affect the inspiral, and the detectability of these effects by LISA Cardoso and Maselli (2020); Kavanagh et al. (2020).
@@ X GW probes of atomic DM
It is possible that the dark sector is actually much more complicated than a single new particle with a single coupling to gravity or the standard model. In fact, atomic DM models exist Feng et al. (2009); Kaplan et al. (2010) in which dark protons and dark electrons interact via a massless dark photon [^12] This dark photon is different from both the dark photon that couples to baryons, discussed in Section II.4, and the one that kinetically mixes with the ordinary photon, discussed in Section VI.3.3. with a particular coupling. In such a model, DM can form bounds states analogous to the Hydrogen atom, and dissipate energy via Bremsstrahlung, recombination, and collisions Buckley and DiFranzo (2018); Rosenberg and Fan (2017). If this kind of DM collects in small-enough regions of space, it will collapse and form “dark BHs” Shandera et al. (2018). Such dark BHs could form in binary systems, as astrophysical BHs do; thus, LIGO, Virgo and KAGRA observations of binary black hole mergers can constrain the existence of these objects, just as they do for PBHs Singh et al. (2021); Abbott et al. (2022f, 2023).
Here, the observable GW signature of dark BHs would come from the inspiral, merger and ringdown of binary black hole systems. However, this signature is not unique: there does not yet exist a way of distinguishing between astrophysical, dark BHs and PBHs. Thus, the constraints that we present in Section X.1 and Section X.2 on BHs with $M_{\rm BH}=[1,100]M_{\odot}$ and $M_{\rm BH}=[0.1,1]M_{\odot}$, respectively, assume that the BHs that are searched for in LIGO, Virgo and KAGRA data are dark BHs.
X.1 Constraints from observations of binary black holes
Ref. Shandera et al. (2018) first considered the prospects of detecting dark BH mergers with advanced LIGO and ET, finding that these interferometers could set stringent constraints on the fraction of DM that dark BHs could compose. If the BH merger of GW190425 Abbott et al. (2020a) is interpreted as the merger of two dark BHs, the Chandrasekhar mass Chandrasekhar (1931) can be constrained to be below $1.4M_{\odot}$ at $99.9\%$, bounding the lower-most value mass of the dark proton as 0.95 GeV and the spacing of energy levels in a dark atom to be near $10^{-3}$ eV Singh et al. (2023). In order to obtain these constraints, we note that LIGO, Virgo and KAGRA observations of binary black holes provide estimates of the chirp mass and the component masses of the system. Ref. Singh et al. (2021, 2023) jointly constrain, in a Bayesian way, the fraction of DM that dark BHs could compose and the minimum mass $M_{\rm min}$ of the distribution of dark BHs assumed, while marginalizing over other hyperparmeters, such as the slope and range of the mass distribution. These posteriors are conditioned on the output of a GW search, which is essentially predicted merger rate density that depends both on what binary black holes have been observed and power spectral density of the interferometers.
Quantitatively, these constraints on the mass of the dark proton and the Chandrasekhar limit of the mass of dark BHs are derived using Chandrasekhar (1931); Shandera et al. (2018); Singh et al. (2021):
@@ X.1 Constraints from observations of binary black holes
where $m_{p}$ is the mass of the proton. The range of the dark proton mass $m_{\chi}$ can be obtained by setting $M_{\rm DC}=M_{\rm min}$ for the values of $M_{\rm min}$ for which the fraction of DM that dark BHs could compose is less than 1:
$\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm max}}}m_{p}<m_{\chi}<\sqrt{\frac{1.4M_{\odot}}{M_{\rm min}^{\rm min}}}m_{p}$ (113)
In Singh et al. (2021), $M_{\rm min}=[0.054,1.50]M_{\odot}$, which implies $0.91<m_{\chi}/\,\mathrm{GeV}<4.76$ when considering both GW190425 and GW190814 Abbott et al. (2020b) as dark binary black hole mergers.
@@
Beyond a range of implied $M_{\rm min}$ and $m_{\chi}$, constraints can be placed on the fraction of DM taht PBHs compose as a function of the average mass of the binary black hole system. These constraints are shown in Fig. 27, assuming that (1) none of the detected BHs by LIGO, Virgo and KAGRA are dark, (2) half are dark, and (3) GW190425 and/or GW190814 are dark. Here, we limit ourselves to case (1) and case (3). These constraints, along with those on the dark proton mass, rely on an assumed mass distribution of dark BHs based on Population III star formation works Bromm and Larson (2004), which is chosen to be $P(m)\propto m^{-b}$ along with a mass range of $M_{\rm min},rM_{\rm min}]$. The parameters b=[1,2],r=[2,100]b=[-1,2],r=[2,100] determine the mass distribution of dark BHs. Assuming all BHs are not dark, the constraint beats existing ones by several orders of magnitude; assuming that GW190425 and GW190814 are dark BHs, a region in this parameter space is carved out of possible average mass/ DM fractions. More observations of dark BHs would be needed to further narrow this region. However, it must be emphasized that while GW190425 and GW190814 being dark BHs are not yet excluded by other experiments, this work does not provide a way of distinguishing astrophysical BHs from dark BHs.
@@
Figure 27: Taken from Singh et al. (2021). Constraints on the fraction of DM that dark BHs could compose relative to other experimental upper limits using all binary black holes detected in the first three observing runs of LIGO, Virgo and KAGRA. These plots assume an atomic model of DM that forms dark BHs. The constraints are obtained in three cases: (1) assuming none of the detected BHs are dark; (2) assuming only GW190425 is a dark BH; and (3) assuming both GW190425 and GW190814 are dark BHs. The latter two cases essentially arise from a joint posterior distribution conditioned on the rate densities inferred from the search results, which also constrain the minimum allowed mass in the assumed population distribution (see text for details). Even in the absence of any dark BHs, the constraints obtained in Singh et al. (2021) significantly outperform existing microlensing experiments Allsman et al. (2001); Tisserand et al. (2007); Niikura et al. (2019) and those from dynamics of dwarf galaxies Brandt (2016); Koushiappas and Loeb (2017). Additional constraints arise from PBH searches Abbott et al. (2019); Gaggero et al. (2017); Raidal et al. (2017, 2019); Ali-Haimoud et al. (2017). SB stands for stochastic background and refers to analyses of the first and second observing runs of LIGO, Virgo and KAGRA data for stochastic GW backgrounds from generic systems.
@@
X.2 Constraints from sub-solar mass binary black hole searches
@@
The lack of smoking-gun evidence for dark BHs in current LIGO, Virgo and KAGRA observations motivates the need to consider sub-solar mass BHs, whose origins, if detected, cannot be explained by astrophysical means. Searches for GWs from sub-solar mass compact objects in binary systems have grown in popularity in recent years, covering mass ranges of roughly $10^{-11},1]M_{\odot}$ Nitz and Wang (2021); Phukon et al. (2021); Abbott et al. (2022f); Miller et al. (2022a); Nitz and Wang (2022); Abbott et al. (2023, 2022b); Miller et al. (2024), and, traditionally, have put constraints on the fraction of DM that PBHs could compose. However, as discussed in Singh et al. (2021), these searches do not assume a particular formation mechanism for these compact objects, and the binaries can thus be interpreted as arising from not only PBHs, but also dark BHs.
@@
In the most recent analysis of LIGO, Virgo and KAGRA data from the second half of the third observing run, constraints on dark BHs with masses below $1M_{\odot}$ were placed, as shown in Fig. 28, using different matched filtering methods, again assuming a power-law dependence for the mass distribution of dark BHs and as a function of the minimum mass of this distribution. These limits do not go below $\sim 10^{-2}M_{\odot}$ because the maximum mass probed is between $2,1000]M_{\odot}$, and so the distribution tends to disfavor smaller masses. Moreover, the limits do not go above a solar mass because the search is only for sub-solar mass objects.
@@
From this work Abbott et al. (2023), the mass of the dark proton has been constrained to be [0.66,8.8] GeV, which is actually broader than the constraint in Singh et al. (2021) of [0.91,4.76] GeV since there are no assumed detections of dark BHs.
Additionally, constraints can be placed on the energy spacing between dark atomic levels. In atomic DM models, the collapse of dark gas into compact stars (and hence into dark BHs) requires efficient radiative cooling, which is governed by the energy level spacing
$\Delta E\sim\tfrac{1}{2}\,\alpha_{\chi}^{2}m_{e_{\chi}},$ (114)
where $\alpha_{\chi}$ is the dark fine-structure constant and $m_{e_{\chi}}$ the mass of the dark electron. By combining the Chandrasekhar mass relation with assumptions about the dark fine-structure constant, the dark fermion masses, and the Jeans mass for collapse, one can use Eq. (1) of Singh et al. (2021) together with Eq. (4) of Shandera et al. (2018) to infer the allowed range of $\Delta E$. For the GW190425 interpretation, this yields a characteristic value of order $\sim 10^{-3}\,\mathrm{eV}$.
@@
Figure 28: Taken from Abbott et al. (2023). Constraints on the fraction of DM that dark BHs could compose in the atomic DM scenario. A mass distribution for dark BHs is assumed with the shown minimum mass on the $x$-axis, and the constraints are shown for three different matched filtering algorithms: GstLAL (dotted), MBTA (dashed) and PyCBC (solid). Constraints from a previous search for sub-solar mass compact objects in the first half of the third observing run of LIGO, Virgo and KAGRA are also shown  Abbott et al. (2022f). This plot differs from Fig. 27 primarily in the mass range that is considered in the sub-solar mass search versus the ones in, e.g. Abbott et al. (2021c). At lower masses, the constraints on $f_{\rm DBH}$ weaken with respect to those at higher masses.
Although sub-solar mass searches for dark BHs have placed promising constraints, they are typically limited to masses above $\sim 10^{-2}M_{\odot}$, largely due to the prohibitive computational cost of conducting matched filtering searches for lighter-mass binaries. In contrast, CW methods—which search for time-frequency tracks rather than relying on matched filtering —can access lower-mass regimes Miller et al. (2021b); Abbott et al. (2022b); Miller (2024b); Alestas et al. (2024). While originally developed for PBH searches, these techniques are equally applicable to testing the dark BH hypothesis. As such, CW approaches offer a valuable path toward exploring a broader parameter space than is feasible with traditional matched filtering methods.